What is the correct equation for the following:
Group goes to show, total ticket price $720, 4 people drop out, each ticket goes up $2. What is new ticket price?
My equation was 720/x + 720/(x-4) = 2, but I don't think that is correct!
Thanks!

Found 2 solutions by edjones, rapaljer:
Answer by edjones(8007)

Let n=number of people and x=ticket price.
nx=720
(n-4)(x+2)=720 group price must remain the same.
nx+2n-4x-8=720
n=720/x
720x/x + 2*720/x - 4x - 8 = 720
720 + 1440/x - 4x - 8 = 720
1440/x - 4x - 8 = 0
1440-4x^2-8x=0
4x^2+8x-1440=0
x^2+2x-360=0
(x-18)(x+20)=0
x=18
18+2=$20.00 new ticket price
.
Ed

Answer by rapaljer(4671)

Let x=original ticket price
720/x= number of tickets originally

Equation is based upon NEW ticket price times NEW number of tickets = 720
(x+2)(720/x - 4)=720
720 - 4x +1440/x - 8 = 720
-4x +1440/x - 8 = 0
-4x^2+1440 -8x=0
x^2 +2x -360=0
(x+20)(x-18)=0
x=-20 reject
x=$18 = Original Ticket Price
x+2= $20 = New Ticket Price

Check:
Number of tickets originally =40
Number of tickets after dropouts = 36

18*40 = 20*36 = $720

Nice Problem!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida